OpenStudy (anonymous):

evaluate the definate integral 1. integrate sec^2t square root tant dt from 0 to pi/4

OpenStudy (rogue):

\[I = \int\limits_{0}^{\pi/4} \sec^2 (t) \sqrt { \tan (t) } dt\]I'll give you a hint. Solve via substitution; Let u = tan (t).